Coronado Industries produces flash drives for computers, which it sells for $25 each. Each flash drive requires $6 of variable costs to make. During April, 1,000 drives were sold. Fixed costs for April were $1,000. How much is the contribution margin ratio?
Answer #1

Contribution margin per unit = sales - variable cost

= 25 - 6 = $19

Contribution margin ratio = Contribution margin / sales

= 19/25 = 76%